OpenStudy (anonymous):

what is the LCM of 35z^4 and 245z^6? Help

OpenStudy (anonymous):

245z^6

OpenStudy (anonymous):

the LCM is \[245z ^{6}\] The GCF is \[35z ^{4}\]

OpenStudy (anonymous):

because \[245z^6=245z^6\times 1\] and \[245z^6=35z^4\times 7z^2\]

OpenStudy (anonymous):

asked for least common multiple yes? not greatest common factor

OpenStudy (anonymous):

yes i asked for the least common multiply

OpenStudy (anonymous):

so what does that make the LCM then please?